# Unifying Multitrack Music Arrangement via Reconstruction Fine-Tuning and Efficient Tokenization

### v2

@@ 3.3.2 Piano Reduction
This task focuses on simplifying multi-instrumental musical pieces into solo piano arrangements that ensure pianistic playability while preserving the original musical essence, i.e., harmonies and textures. In the training process, $\operatorname{\mathcal{T}_{\text{task}}}$ is defined to extract the original piano tracks, while $\operatorname{\mathcal{S}_{\text{task}}}$ is an identity transformation. To ensure the piano part is sufficiently prominent in the training data, we filter the dataset to include only segments where the piano pitch range constitutes more than 40% of the total pitch range, reducing the dataset to approximately 40% of its original size. The segment length is set to 1 bar. Additionally, drum tracks are also pre-removed from $y$.
3.3.3 Drum Arrangement
This task involves creating a drum track for songs that lack one. The model needs to recognize the groove of the music and enhance it using the drum set, and further, handle transitions between musical phrases to drive the music forward and make it more engaging, which consequently requires a better understanding of musical structure. Here, $\operatorname{\mathcal{S}_{\text{task}}}$ extracts tracks belonging to pitched instruments, while $\operatorname{\mathcal{T}_{\text{task}}}$ extracts the drum track. A longer segment length of 4 bars is used for this task, as drum patterns generally span across multiple bars.
@@ 4.1 Implementation Details
Our model, an 80M-parameter decoder-only Transformer, has a hidden dimension of 768, 12 layers, 16-head attention, and a context length of 2048 tokens (around 8$\times$ the longest bar in our dataset). The model first undergoes a standard LM pre-training, and then was fine-tuned with the proposed objective. Pre-training used four RTX A5000 GPUs (batch size 12, 1 epoch), while fine-tuning used a single A40 GPU (variable batch size, 3 epochs). Training was done with two publicly accessible datasets. Pre-training adopted the Los-Angeles-Project dataset Lev (2024) (405K MIDI files, 4.3B tokens after REMI-z tokenization, 2% validation split) and fine-tuning was done with Slakh2100 Manilow et al. (2019) (1,289 training, 270 validation, 151 test MIDI files), featuring 34 pitched instruments and drums, with $\geq$5 tracks per piece. Detailed hyperparameter settings are in Appendix B.
4.2 Tokenizer Efficiency Comparison
We compare REMI+ and REMI-z through theoretical and empirical analyses. First, we tokenize the Slakh2100 dataset using both methods and evaluate compactness via the average number of tokens per bar ($\bar{T}_{\text{bar}}$) and per note ($\bar{T}_{\text{note}}$). Lower values indicate reduced computational costs during training and autoregressive decoding.
Next, we quantify token sequence complexity using bar-level Shannon entropy. For a discrete random variable $X$ representing token distributions, entropy of a token sequence $H(X)$ is defined as:
@@ 4.2 Tokenizer Efficiency Comparison
where $\{x_{1},\dots,x_{N}\}$ are all tokens appeared in the sequence, and $\{P(x_{1}),\dots,P(x_{N})\}$ are their probabilities. We calculate the entropy for each bar sequence to get $H_{\text{bar}}$, and then average across all bars in the dataset to obtain $\bar{H}_{\text{bar}}$. Lower $\bar{H}_{\text{bar}}$ value indicates less inherent complexity of the bar sequence, and by extension, the multitrack music, reflecting efficient encoding with less redundant information.
Then, we train unconditional LMs on the Slakh2100 dataset and observe the note-level perplexity:
where $M$ is the number of notes in a bar, and $P(n_{j}\mid n_{1:j-1})$ aggregates probabilities for all tokens corresponding to note $n_{j}$ (instrument, position, pitch, duration). Unlike token-level perplexity ($\text{PPL}_{\text{token}}$), this metric normalizes by the number of notes, ensuring fair comparison across tokenization schemes. Lower note-level perplexity indicates better modeling of musical notes, as it reflects reduced uncertainty in predicting all associated attributes.
Finally, we assess task-specific performance with different tokenizers, discussed in the next section.
4.3 Task-Specific Evaluation
This section introduces baseline models and metrics used for evaluating the music arrangement models.
4.3.1 Baseline Models
@@
For each task, we compare our model against a state-of-the-art (SOTA) task-specific baseline. For band arrangement, we adopt Transformer-VAE from Zhao et al. (2024), the strongest previously reported model for multitrack arrangement without assumptions on track type or number. It combines Transformer-based long-term and inter-track modeling with a VQ-VAE generation module. For piano reduction, we compare with Terao et al. (2023) (UNet), the most recent work in this area. For drum arrangement, existing drum-specific models (e.g., Barnabò et al. (2023), Dahale (2022)) are unsuitable as they generate drums from a single melody or instrumental track. Instead, we adopt Composer’s Assistant 2 (CA v2) Malandro (2024), a strong track infilling model capable of handling multitrack inputs and generating drum outputs. Baseline’s implementation details are in Appendix B.
@@
To demonstrate the impact of generative pre-training, an ablation variant of our model without pre-training (w/o PT) was used as a baseline. Additionally, we include rule-based baselines to showcase the difficulty of the tasks. For band arrangement, Rule-Based distributes notes evenly by pitch across instruments. For piano reduction, we use Rule-F (flattened multitrack where the piano plays all notes) and Rule-O (original piano track). Note that Rule-O is not considered ground truth since the piano track is likely to not fulfil the $\geq 40\%$ pitch range requirements. For drum arrangement, the original drum track (Ground Truth) is included anonymously for human evaluation as a golden standard.
4.3.2 Segment-level Objective Evaluation
@@
Objective metrics measure similarity between model outputs and target sequences, assuming closer resemblance indicates higher naturalness and musicality. Given the extreme sparsity of note events in the track-wise piano roll (e.g., only 0.04% non-zero elements in the Slakh2100 dataset under 48th-note quantization), F1-based metrics are more suitable than accuracy-based metrics for similarity evaluation. Following Malandro (2024); Terao et al. (2023), we use note-level F1 to measure similarity between model outputs and target sequences. Specifically, we compute Note F1 (correct onset and pitch) and Notei F1 (additional correct instrument prediction), both under 16th-note quantization for fair comparison.
@@
For piano reduction and drum arrangement, the same models are used in objective and subjective evaluations. For band arrangement, models are separately trained without random track deletion to ensure deterministic outputs. Baseline models are also modified by excluding the prior model to remove long-term context hints, ensuring evaluation fairness.
@@
In addition to the modifications described above, we introduce three additional metrics specific to the band arrangement task. First, Instrument Intersection over Union (I-IoU) evaluates the accuracy of instrument control. Second, Voice Word Error Rate (V-WER) measures the similarity in voice features between the generated output and the reference, indicating the effectiveness of the voice control. Third, Melody F1 (Mel F1) calculates Note F1 on the tracks with the highest average pitch in the output and target music, assessing the preservation of melody, which is critical for arrangement fidelity. The detailed computation of these metrics is provided in Appendix C.
Beyond comparing with the SOTA model, the objective evaluation for band arrangement serves two additional purposes: (1) to assess the task-specific performance of different tokenizers and (2) to validate the design of the fine-tuning objective. The band arrangement task is particularly challenging because it imposes no assumptions on the types or numbers of target instruments, making it more representative of the model’s overall capabilities compared to the other tasks.
4.3.3 Song-Level Subjective Evaluation
@@
To complement similarity metrics and account for perceptual quality and creativity, we conducted human evaluations. Full-piece arrangements were generated by all models and evaluated phrase-by-phrase on a 5-point scale (1: very low, 5: very high). For band arrangement, models were tested across three instrument combinations with different complexity: string trio (3 tracks), rock band (4 tracks), and jazz band (7 tracks). Three metrics were used across band, piano, and drum arrangement tasks: Coherence, which evaluates the natural flow of the arrangement and the consistency of each instrument’s performance and style throughout the piece; Creativity, which assesses the degree of innovation in the arrangement under the constraints of the music’s content and style; and Musicality, which measures the overall musical appeal and aesthetic quality of the arrangement. Further details on the metrics, ensemble settings, questionnaire, and evaluation process are in Appendix D.
@@
Task-specific metrics were also introduced. For band arrangement, Faithfulness measures resemblance to the original in melody and overall feel, while Instrumentation assesses the appropriateness of each instrument’s role within the band and their harmony. For piano reduction, Faithfulness is also adopted but without melody preservation requirements, and Playability assesses the feasibility of the generated contents played by human pianists. For drum arrangement, Compatibility measures how well the drum track blends with other instruments, and Phrase Transition assesses the smoothness of transitions between phrases.
Additionally, significance tests were conducted between our model and the SOTA baselines using within-subject (repeated-measures) ANOVA Scheffe (1999). The significance level are showed on the result tables: * for p-values less than 0.05, $\dagger$ for p-values less than 0.01, and $\ddagger$ for p-values less than 0.001.
@@ 5 Results
5.1 Effectiveness of REMI-z
@@
Table 2: Tokenizer efficiency comparison.
As shown in Table 2, our proposed tokenization method significantly reduces the average sequence length from 225.91 tokens (REMI+) to 151.68 tokens per bar, achieving a 32.9% reduction. The number of tokens per note also decreases from 4.03 (REMI+) to 2.77 with REMI-z. These results demonstrate that our tokenizer generates a more compact representation for multitrack music, which not only reduces the computational complexity for generative modeling but also accelerates autoregressive decoding and potentially facilitates the modeling of long-term dependencies.
Moreover, this increased compactness does not introduce additional complexity. In terms of bar-level Shannon entropy, REMI-z achieves 29.43 bits/token compared to REMI+’s 41.68 bits/token, representing a 29.4% relative reduction. This indicates that sequences tokenized with REMI-z contain less redundant information and can be further encoded with a shorter expected code length, requiring simpler encoding rules for lossless compression and potentially enabling more effective generative modeling.
We confirm this advantage with empirical analysis. The LM trained with REMI-z achieves lower note-level perplexity (84.11 vs 116.20), demonstrating that modeling the basic unit of music—the note—is more effective with REMI-z. This improvement in note-level modeling naturally extends to better performance in generating bars and complete compositions. Conversely, while individual REMI+ tokens are easire to model (-1.5 token-level perplexity), this does not translate to better note-level and song-level performance.
5.2 Band Arrangement
Model I-IOU V-WER Note F1 Notei F1 Mel F1
Transformer-VAE 97.5 35.0 49.5 40.0 24.7
Transformer w/ REMI+ 95.0 18.2 94.4 76.0 68.8
Transformer w/ REMI-z 99.5 9.9 97.8 77.5 77.8
$+$ Pre-training (Ours) 99.8 7.6 97.5 87.0 84.5
-- voice 99.6 17.6 97.2 84.3 81.5
-- history 100.0 9.0 97.6 77.4 79.4
Table 3: Objective evaluation results of the band arrangement task.
5.2.1 Objective Evaluation
We then evaluate REMI-z’s task-specific performance. As shown in Table 3, REMI-z consistently outperforms REMI+ on all metrics. It achieves better instrument control (-4.5% I-IOU), reducing undesired or missing instruments, and more accurate voice control with lower V-WER (-8.3%). This indicates REMI-z enable the model to better follows user-specified instrument constraints. The higher Note F1 and Notei F1 scores demonstrate improved reconstruction quality, highlighting REMI-z’s effectiveness in the arrangement training. The significantly higher Mel F1 (+9.0%) also indicates better melody preservation and arrangement fidelity.
Pre-training further enhances our model’s performance in three major aspects: (1) improved voice control (-2.3% V-WER), indicating better understanding of voice requirement in the instrumental prompts, (2) more accurate per-instrument reconstruction (+9.5% Notei F1), suggesting enhanced comprehension of instrumental roles in ensembles, and (3) better melody preservation (+6.7% Mel F1), leading to improved arrangement fidelity. For detailed knowledge probing analysis of how pre-training aligns learned representations with musical properties, please refer to Appendix E.
Our model consistently outperforms Transformer-VAE across all metrics. The gap is particularly significant in Note F1 (97.5% vs 49.5%) and Notei F1 (87.0% vs 40.0%), indicating Transformer-VAE’s limitation in inferring instrument-specific content within musical contexts. Our model also achieves substantially better melody preservation (+59.8% Mel F1), ensuring higher-fidelity song-level arrangements.
Ablation studies validate our design choices for voice and history conditioning. Excluding relative voice information from the instrument condition sequence leads to degraded voice relationships (-10.0% V-WER) and consequently lower Notei F1 (-2.7%) and Mel F1 (-3.0%). This demonstrates that voice conditioning enables effective voice control and facilitates instrument role inference. Similarly, excluding history from the condition sequence noticeably reduces Notei F1 (-9.6%) and Mel F1 (-5.1%), laying the foundation for generating coherent outputs in song-level inference.
5.2.2 Subjective Evaluation
Model Fa. Co. In. Cr. Mu.
Rule-Based 3.46 3.05 2.89 3.00 3.07
Transformer-VAE 2.65 2.70 2.72 3.00 2.72
Ours 3.77 3.47 3.49 3.40 3.47
w/o PT 3.19 2.82 2.86 2.93 2.75
As shown in Table 4, our model consistently achieves the highest scores across all metrics, demonstrating its capability to generate high-quality arrangements that adhere to the original musical essence while achieving good Coherence, Creativity, and Musicality. Transformer-VAE scores lower than the Rule-Based model, particularly in Faithfulness and Musicality, indicating its limitations in preserving musical essence and expression. This also suggests that the rule-based method can serve as competitive baselines for band arrangement research. Notably, removing pre-training degrades all metrics, highlighting its importance for generation quality.
5.3 Piano Reduction
Model F1 Fa. Co. Pl. Cr. Mu.
Rule-F - 3.93 3.59 3.14 2.96 3.34
Rule-O - 2.75 3.49 4.07 2.62 2.96
UNet 58.3 2.97 2.90 3.47 2.82 2.78
Ours 85.5 3.63 3.64 3.86 3.14 3.48
w/o PT 78.4 2.25 2.58 3.29 2.67 2.26
Table 5 presents the piano reduction results. Our model outperforms all baselines in F1 score, Coherence, and Musicality. Unlike Rule-O and Rule-F which excel only at Playability or Faithfulness respectively, our model achieves a balance between these two aspects. Pre-training also proves essential: without it, the model has lower F1 and underperforms UNet from Terao et al. (2023) across all subjective metrics. In contrast, the model with pre-training significantly surpasses UNet in subjective evaluations. Interestingly, F1 scores do not strictly correlate with perceptual quality, as evidenced by the w/o PT model’s higher F1 but lower subjective evaluation scores compared to UNet.
5.4 Drum Arrangement
Model F1 Comp. Co. Tr. Cr. Mu.
Ground Truth 100.0 4.31 4.18 3.36 3.16 3.78
CA v2 20.3 3.82 4.05 2.86 2.58 3.19
Ours 79.3 3.91 4.03 3.77 3.27 3.57
w/o PT 1.2 2.49 2.19 2.21 2.82 2.05
As in Table 6, our model significantly outperforms baseline CA v2 in F1 score. On subjective metrics, our model performs comparably with CA v2 on Coherence, and better at all remaining metrics. Particularly, we notice that CA v2 tends to generate repetitive content throughout the song. Conversely, our model excels in handling phrase transitions and demonstrating creativity, achieving a Musicality score comparable to the ground truth. This indicates that our model effectively captures high-level musical structures during generation, introducing variations between conventional drum patterns and producing more engaging arrangements. The model without pre-training fails to perform meaningfully (1.2% F1), with a notably larger performance gap compared to other arrangement tasks. This demonstrates pre-training’s crucial role in sequence-to-sequence tasks when handling longer musical structures, where complex alignments between source and target sequences require enhanced structural understanding.

### v3

@@ 3.3 Task Instantiations
Piano Reduction. This task simplifies ensemble music into solo piano accompaniment, aiming to preserve key harmonic and textural elements while ensuring pianistic playability. The $\operatorname{\mathcal{T}_{\text{task}}}$ selects piano tracks, while $\operatorname{\mathcal{S}_{\text{task}}}$ is the identity mapping. To ensure the reduction is meaningful, we filter training data to keep only segments where the piano part is sufficiently prominent (covering >>40% of the pitch range). Drum tracks are removed from input, and the segment length is set to 1 bar.
Drum Arrangement. The goal of this task is to generate a drum track for music that lacks one. Here, $\operatorname{\mathcal{S}_{\text{task}}}$ extracts all pitched-instrument tracks, while $\operatorname{\mathcal{T}_{\text{task}}}$ extracts the drum track. We use 4-bar segments since drum patterns sometimes span multiple bars. The model must recognize the underlying groove and phrase boundaries of the source music, enhance them with coherent and stylistically appropriate rhythmic patterns, and ensure proper transitions across segments—requiring a stronger understanding of rhythmic and structural organization.
@@ 4.1 Implementation Details
4.2 Baseline Models
@@
For each task, we compare our model against a state-of-the-art (SOTA) task-specific baseline. For band arrangement, we adopt Transformer-VAE from zhao2023accomontage , the strongest previously reported model for multitrack arrangement without assumptions on track type or number. It combines Transformer-based long-term and inter-track modeling with a VQ-VAE generation module. For piano reduction, we compare with terao2023neural (UNet), the most recent work in this area. For drum arrangement, we adopt Composer’s Assistant 2 (CA v2) ComposersAssistant2 , a SOTA track infilling model capable of handling multitrack inputs and generating drum outputs. Existing drum-specific models (e.g., barnabo2023cycledrums , dahale2022automatic ) are unsuitable for our setting, as they assume a single melody or instrumental track input rather than general multitrack conditioning. Baseline’s implementation details are in Appendix B.4.
@@
To demonstrate the impact of generative pre-training, an ablation variant of our model without pre-training (w/o PT) is used as a baseline. Additionally, we include simple rule-based baselines as non-learning references, designed to provide naive solutions with minimal musical knowledge, helping to contextualize the difficulty of arrangement tasks. For band arrangement, Rule-Based distributes notes evenly by pitch across instruments, serving as a naive reinterpretation strategy. For piano reduction, we use Rule-F (a flattened multitrack where the piano plays all notes), which reflects an overcomplete reduction prioritizing coverage, and Rule-O (the original piano track), which provides a playability-guaranteed but musically incomplete reduction. For drum arrangement, the original drum track (Ground Truth) is included anonymously in the human evaluation as an upper bound on perceptual scores.
4.3 Objective Evaluation
@@
Objective metrics measure similarity between model outputs and target sequences, assuming closer resemblance to human-created music indicates higher naturalness and musicality. Following ComposersAssistant2 ; terao2023neural , we use note-level F1 to measure similarity between model outputs and target sequences. Specifically, we compute Note F1 (correct onset and pitch) and Notei F1 (additional correct instrument prediction), both under 16th-note quantization for fair comparison with baselines.
@@
For piano reduction and drum arrangement, the same models are used in objective and subjective evaluations. For band arrangement, models are separately trained without random track deletion to ensure deterministic outputs. Baseline models are also modified by excluding its prior model to remove long-term context hints, ensuring evaluation fairness.
@@
In addition to the modifications described above, we introduce three task-specific metrics for band arrangement. First, Instrument Intersection over Union (I-IoU) evaluates the accuracy of instrument control. Second, Voice Word Error Rate (V-WER) measures the similarity in voice features between the generated output and the reference, reflecting how well the model follows the voice conditions specified by instrument token ordering. Third, Melody F1 (Mel F1) computes the Note F1 score on melody tracks, estimated as the tracks with the highest average pitch in the output and reference, to assess how well the original melody is preserved—an important factor for perceived fidelity. Detailed definitions and computation procedures are provided in Appendix C.1.
Among all tasks, band arrangement is a strong testbed for evaluating controllability and generalizability because it requires the highest flexibility without assumptions on target instrument types or counts. Hence it serves two additional purposes: (1) to validate key design choices in our fine-tuning objective, particularly the use of voice-aware instrumentation and segment-level history conditioning; and (2) to prove the effectiveness of the proposed tokenization schemes in arrangement task. When comparing tokenization schemes, statistical significance is computed by Wilcoxon signed rank test Wilcoxon1945IndividualCB . For a broader analysis of tokenization’s impact on unconditional generation, see § 5.4.
4.4 Human Evaluation
@@
To complement similarity-based objective metrics and assess perceptual quality and creativity, we conducted human evaluations. Full-piece arrangements were generated by all models and evaluated on a 5-point scale (1: very low, 5: very high). For band arrangement, models were tested across three instrument combinations with different complexity: string trio (3 tracks), rock band (4 tracks), and jazz band (7 tracks). Three metrics were used across band, piano, and drum arrangement tasks: Coherence, which evaluates the natural flow of the arrangement and the consistency of each instrument’s playing style throughout the piece; Creativity, which assesses the degree of innovation in the arrangement under the constraints of the music’s content and style; and Musicality, which measures the overall musical appeal and aesthetic quality of the arrangement. Further details on the metrics, ensemble settings, questionnaire, and evaluation process are in Appendix D.
@@
Task-specific metrics were introduced for the distinct evaluation needs of each arrangement scenario. For band arrangement, Faithfulness measures resemblance to the original in melody and overall feel, while Instrumentation assesses the appropriateness of each instrument’s role within the ensemble and their harmony. For piano reduction, Faithfulness is also adopted but without melody preservation requirements, and Playability assesses the feasibility of the generated contents played by human pianists. For drum arrangement, Compatibility measures how well the drum track blends with other instruments, and Phrase Transition assesses the smoothness of transitions between musical phrases. We report mean and standard error of mean in result tables. Significance tests were conducted between our model and the SOTA baselines using within-subject (repeated-measures) ANOVA scheffe1999analysis .
@@ 5 Results
5.1 Band Arrangement
5.1.1 Objective Evaluation
Table 1: Objective evaluation results for the band arrangement task. Statistical significance is indicated as follows: * for $p<0.05$, $\dagger$ for $p<0.01$, and $\ddagger$ for $p<0.001$.
Model I-IOU ↑ V-WER ↓ Note F1 ↑ Notei F1 ↑ Mel F1 ↑
Transformer-VAE 97.5 35.0 49.5 40.0 24.7
Transformer w/ REMI+ 95.0 18.2 94.4 76.0 68.8
Transformer w/ REMI-z 99.5 9.9 97.8 77.5 77.8
++ Pre-training (Ours) 99.8 7.6 97.5 87.0 84.5
- voice 99.6 17.6 97.2 84.3 81.5
- history 100.0 9.0 97.6 77.4 79.4
Tokenization impact. Table 1 shows that model adopted REMI-z significantly outperforms that adopted REMI+ across all objective metrics. It enables stronger instrument control (I-IOU: 99.5% vs. 95.0%) and voice control (V-WER: 9.9% vs. 18.2%), aligning outputs better with user-specified conditions. The improvement in Note F1 (+6.7%) confirms higher reconstruction quality, while the gain in Notei F1 (+1.9%) highlights enhanced instrument-wise modeling. Furthermore, the substantial increase in Mel F1 (+9.0%) suggests REMI-z enable the model to better identify the melody components from content sequence, which is improtant for arrangement fidelity.
Pre-training benefit. Pre-training brings clear gains by transferring musical knowledge useful for arrangement. It improves three key aspects: (1) lower V-WER ($2.3\%$), indicating more effective voice control; (2) higher Notei F1 ($9.5\%$), reflecting enhanced instrument-wise content modeling; and (3) the highest Mel F1 (84.5%), indicating stronger melody preservation. Probing analysis (Appendix E) further shows that pre-training strengthens the alignment between token embeddings and musical concepts useful for arrangement, such as instrument roles and chord progression.
Comparison with Transformer-VAE. Our model outperforms Transformer-VAE by a wide margin. In particular, it achieves much higher Note F1 (97.5% vs. 49.5%), Notei F1 (87.0% vs. 40.0%), and Mel F1 (84.5% vs. 24.7%), highlighting the advantage of our context-aware, content- and instrument-conditioned generation approach over the latent inference used in Transformer-VAE.
Ablation: voice and history conditioning. Removing voice information from instrument conditions degrades voice control (V-WER: $10.0\%$), demonstrating the effectiveness of our voice control method. It also lowers per-instrument F1 (Notei: $2.7\%$) and Mel F1 ($3.0\%$), indicating that voice-order information serves as a useful hint for inferring instrument roles during arrangement. Excluding history conditioning results in even larger drops in Notei F1 ($9.6\%$) and Mel F1 ($5.1\%$), validating that temporal context facilitates accurate reconstruction, laying the foundation for coherent song-level arrangement.
5.1.2 Human Evaluation
Model Fa. ↑ Co. ↑ In. ↑ Cr. ↑ Mu. ↑
[…]
Strong subjective performance. As shown in Table 2, Our model achieves the highest scores across all subjective metrics than all baselines, demonstrating its ability to generate coherent, stylistically appropriate, and musically appealing arrangements while preserving core musical essence.
Transformer-VAE underperforms. It lags significantly behind our model in every subjective criterion. Notably, its lower Faithfulness score (-1.12) reflects difficulty in preserving core musical content, while low ratings in Instrumentation (-0.77) and Coherence (-0.77) suggest weaker track-wise modeling and non-idiomatic instrument usage, as well as insufficient coherence between segments. Overall Musicality (-0.75) also falls below our model. It even scores lower than the rule-based baseline, though the latter suffers from inherent limitations in musicality and coherence, making it unsuitable for high-quality arrangement.
Pre-training improves quality. Removing pre-training reduces scores across all metrics, with notable drops in Faithfulness (-0.58) and Musicality (-0.72), reinforcing the importance of musical knowledge transfer from generative pre-training for content retention and overall perceptual quality.
5.2 Piano Reduction
Model F1 ↑ Fa. ↑ Co. ↑ Pl. ↑ Cr. ↑ Mu. ↑
[…]
Best overall quality. As shown in Table 3, our method obtains the highest ratings in F1 (85.5%), Coherence (3.64), Playability (3.86), and Musicality (3.48), while also maintaining strong scores in Faithfulness (3.63) and Creativity (3.14). It significantly outperforms UNet across all subjective metrics, with especially large margins in Faithfulness (+0.66), Coherence (+0.74), and Musicality (+0.70)—indicating better content preservation, cross-segment continuity, and overall musical quality. Pre-training again proves essential: without it (w/o PT), performance drops notably across all metrics.
Balanced fidelity and playability. Compared to rule-based methods, our model achieves a better trade-off between fidelity and playability. Rule-F preserves original content well (Faithfulness: 3.93) but results in poor playability (3.14), while Rule-O achieves the highest playability (4.07) but at the expense of faithfulness (2.75). In contrast, our method produces reductions with balanced Faithfulness (3.63) and Playability (3.86).
5.3 Drum Arrangement
Model F1 ↑ Comp. ↑ Co. ↑ Tr. ↑ Cr. ↑ Mu. ↑
[…]
Best subjective quality among learned models. As shown in Table 4, our model outperforms CA v2 in all subjective metrics except Coherence (4.03 vs. 4.05), and significantly improves Creativity (3.47 vs. 2.58) and Phrase Transition (3.27 vs. 2.86), which are essential for engaging, structurally-aware drum arrangements. It also achieves the highest Musicality score (3.57), closely approaching ground truth (3.78). Again, the w/o PT variant performs poorly across all metrics.
Improved phrasing and variation. Compared to CA v2, which often repeats similar drum patterns throughout a piece, our model produces more varied and context-aware rhythms that better reflect musical phrasing. The higher Phrase Transition score (+0.91) supports this observation, demonstrating the model’s ability to capture structural changes in music during generation.
5.4 Tokenization Efficiency and General Modeling Advantages
Table 5: Tokenization scheme comparison on unconditional generation.
@@ 5.4 Tokenization Efficiency and General Modeling Advantages
Additionally, we evaluate REMI-z against REMI+ to assess its efficiency and effectiveness for generative modeling of symbolic multitrack music. We report several metrics to evaluate compactness of tokenization schemes and their unconditional modeling performance: 1) average tokens per bar ($\bar{T}_{\text{bar}}$), 2) average tokens per note ($\bar{T}_{\text{note}}$), 3) Shannon entropy of bar-level token sequences ($\bar{H}_{\text{bar}}$), and 4) note-level perplexity ($\text{PPL}_{\text{note}}$). Note that $\text{PPL}_{\text{note}}$ is the aggregated conditional probability of all note attribute tokens, normalized by the number of notes, enabling fairer comparison across tokenization schemes. Calculations are detailed in Appendix C.2.
Compactness. Tokenizing Slakh2100 with REMI-z yields a 32.9% reduction in sequence length per bar (151.68 vs 225.91) and fewer tokens per note (2.77 vs 4.03), effectively reducing training and inference computational costs.
Representational Simplicity. REMI-z also produces lower bar-level Shannon entropy (29.43 vs 41.68 bits/token) on Slakh2100 compared to REMI+, despite representing the same underlying musical content. This indicates reduced information redundancy, and suggests that REMI-z sequences consist of more predictable tokens, potentially facilitating the learning of generative models.
Note-Level Modeling Advantage. We train unconditional generation models on Slakh2100 with both tokenizations and compare note-level perplexity with the same architecture as our arrangement models. The model adopting REMI-z achieves substantially lower note-level perplexity (84.11 vs. 116.20), indicating better modeling of complete musical notes. Since notes form the atomic units of music, improvements in note-level modeling directly contribute to lower uncertainty at modeling bars and full compositions, suggesting utility in general symbolic music modeling beyond arrangement tasks. While REMI+ slightly outperforms in token-level perplexity, this does not translate to better modeling of higher-level structures.